# Draft: May 24, 2025

### v1

@@ 5 Compartments
Reacher $c^{2}_{2,2}$ in Fig. 1 Same as in Gymnasium Robotics A2C, ARS, DDPG, PPO, SAC 5.1.1
@@ 5.1 Compartments for Circularity of Solids
The success of sorting $s$ affects the magnitude of $\lambda$. For example, $s$ = 0 yields that all the waste material enters the incinerator at $t_{3,\text{in},6}$, i.e., $\lambda_{\text{a}}$ = $\lambda(\mathcal{N}_{\text{s}};t_{3,\text{in},6},0,T_{\text{s}})$ = $-2m$. In contrast, $s$ = 100 increases circularity by extending the life-cycle of the material, i.e., $\lambda(\mathcal{N}_{\text{s}};t_{3,\text{in},6},100,T_{\text{s}})$ = $-m>\lambda_{\text{a}}$. While all the material is inside the incinerator for $t\geq t_{3,\text{in},5}$ regardless of $s$, i.e.,
@@
the recycled mass $m_{\text{r}}$ has extended the life-cycle of the amount $\nu$ = $t_{3,\text{in},5}$ $t_{3,\text{in},6}$ if $s>0$, which yields an increase of $m_{\text{r}}$ in $\lambda$ for $t_{3,\text{in},6}\leq t<t_{3,\text{in},5}$ ($\lambda$ is independent of $s$ for $t<t_{3,\text{in},6}$ and for $t\geq t_{3,\text{in},5}$). Extending the life-cycle of materials is a key principle of circular economy (Potting et al., 2017).
@@ 5.1.1 Reacher
The Reacher environment (Towers et al., 2024) consists of a robot with two revolute joints rotating around a direction normal to the plane, it has two degrees of freedom, its actions are the torques applied to the revolute joints, and the goal of the robot is to reach a target point generated at a random location. It is a simplification of the pick-and-place task needed to perform waste sorting. Its simplicity enable us to illustrate how the percentage success $s$ is affected by the RL controller performance, which in turns affects $\lambda$ as shown in Fig. 2. We tested the five SB3 algorithms reported in Table 2, where
@@
with $r_{\text{e}}$ and $r_{\text{s}}$ denoting the mean reward over 100 episodes at the end and at the start of the training, respectively. The parameter $s$ is calculated by considering as a success an episode in which the end effector reaches a distance from the target smaller than 0.04 m and the joint torques are less than 0.005 Nm. We have run 10,000 episodes and counted the successful episodes.
Table 2: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$, and $s$ for each algorithm with Reacher. Trainings were executed for 100,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $s$
A2C 04:37 -63.4 -16.6 46.8 17.8
ARS 01:00 -12.1 -11.0 1.1 2.8
DDPG 13:43 -60.2 -5.5 54.7 72.3
PPO 03:59 -60.2 -37.3 22.9 20.6
SAC 22:23 -45.1 -8.3 36.8 51.7
@@ 5.1.2 Fetch Waste Sorting
The Reacher environment is a very simplified model of robotic waste sorting. A more realistic simulator is the one based on the Gymnasium Robotics environment Fetch-PickAndPlace, which is a 7-DOF manipulator that performs pick-and-place with a two-finger gripper (Plappert et al., 2018; Towers et al., 2024). The RL agent controls the opening and closure of the fingers along with the motion of the robot joints. To resemble a waste sorting tasks we modified the Gymnasium environment by constraining the target to be on the table instead of in the air and by locating the target at a fixed point instead of being randomly generated since the coordinates of the point where to place a waste item are usually known and fixed, e.g., the position of the bin. Hence, this environment is referred to as FetchWasteSorting. With this environment, we tested the SB3 algorithms whose performance are reported in Table 3.
Table 3: Training time, $r_{\text{s}}$, $r_{\text{e}}$, and $\zeta$ for each algorithm with FetchWasteSorting. Trainings were executed for 150,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$
DDPG-HER 50:06 -50.25 -50 0.25
SAC-HER 63:43 -50.25 -12.67 37.58
TD3-HER 45:12 -50.25 -50 0.25
TQC-HER 65:19 -50.25 -11.45 38.80
@@ 5.1.3 Transport Truck
The transport truck (compartment $c^{6}_{2,3}$) moves the unsorted waste of mass $m_{\text{u}}$ from the sorting facility ($c^{2}_{2,2}$) to the incinerator ($c^{3}_{3,3}$) as shown in Fig. 1. Currently, the model of the truck in CiRL consists of a particle of mass $m_{\text{tot}}=m_{\text{truck}}+m_{\text{u}}$, where $m_{\text{truck}}$ is the mass of the empty truck. As demonstrated by Zocco et al. (2023), the first law of thermodynamics leads to the Lagrange’s equations, which in turns yield the second Newton’s law of motion (Zocco and Malvezzi, ). Considering that the particle-truck is subject to the traction force $F$, the state-space form of the system is
@@
and hence, the RL controller trained successfully generates a force $F$ that brings the truck to the incinerator located at $x_{\text{inc}}$ and stops it there. This condition corresponds to the maximum reward $r_{\text{max}}=0$. The performance of a single run of the tested RL algorithms is reported in Table 4.
Table 4: Training time, $r_{\text{s}}$, $r_{\text{e}}$, and $\zeta$ for each algorithm with TransportTruck. Trainings were executed for 400,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$
A2C 18:10 $-1.0\times 10^{11}$ $-1.2\times 10^{8}$ $\bm{1.0\times 10^{11}}$
ARS 02:56 $-4.0\times 10^{8}$ $-7.1\times 10^{8}$ $-3.1\times 10^{8}$
DDPG 50:42 $-1.7\times 10^{11}$ $-2.0\times 10^{11}$ $-3.3\times 10^{10}$
PPO 14:34 $-1.03\times 10^{11}$ $-1.02\times 10^{11}$ $\bm{8.6\times 10^{8}}$
SAC 1:28:18 $-9.8\times 10^{10}$ $-1.3\times 10^{11}$ $-3.4\times 10^{10}$
@@ 5.1.4 Incinerator
The compartment $c^{3}_{3,3}$ in Fig. 1 is the incinerator, which processes end-of-life products and materials. From a circular economy perspective, it is among the least circular options because it destroys the functionality of products and their components (Potting et al., 2017). A complete dynamical model of an incinerator was developed by Magnanelli et al. (2020) as a result of mass and energy balances. In this first version of CiRL, we consider only the wastebed and the freeboard sub-compartments of Magnanelli et al. (2020). This yields the following system of six ordinary differential equations in the state-space form and with state vector $\bm{x}=[M,M_{\text{char}},T_{\text{w}},M_{\text{g,w}},M_{\text{g,f}},T_{\text{g}}]^{\top}=[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}]^{\top}$, where $M$ is the mass of waste in the wastebed, $M_{\text{char}}$ is the mass of char in the wastebed, $T_{\text{w}}$ is the temperature in the wastebed, $M_{\text{g,w}}$ is the mass of the gasses in the wastebed, $M_{\text{g,f}}$ is the mass of the gasses in the freeboard, and $T_{\text{g}}$ is the temperature of the gasses in the freeboard.
@@
The performance of a single run of the tested RL algorithms on the incinerator environment is reported in Table 5.
Table 5: Training time, $r_{\text{s}}$, $r_{\text{e}}$, and $\zeta$ for each algorithm with Incinerator. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$
A2C 08:25 $-3.40\times 10^{5}$ $-3.79\times 10^{5}$ $-3.9\times 10^{4}$
ARS 01:28 $-3.65\times 10^{5}$ $-3.53\times 10^{5}$ $\bm{1.21\times 10^{4}}$
DDPG 24:50 $-3.23\times 10^{5}$ $-3.44\times 10^{5}$ $-2.07\times 10^{4}$
PPO 07:30 $-3.51\times 10^{5}$ $-3.58\times 10^{5}$ $-0.71\times 10^{4}$
SAC 43:20 $-3.37\times 10^{5}$ $-3.41\times 10^{5}$ $-0.40\times 10^{4}$
@@ 5.2 Compartments for Circularity of $\text{CO}_{2}$
In this section, we address the circulation of atmospheric carbon dioxide since its accumulation is one of the main causes of climate change (NASA, ). Governments usually refer to the need of carbon dioxide circulation as the “net-zero target”, which is defined by Oxford Net Zero , e.g., the EU (European Commission, a), the UK (UK Government, ), the US (Office of the Federal Chief Sustainability Officer, ), China (International Energy Agency, ), and India (CEEW, ). Using the TMN formalism, the net-zero problem can be represented as
@@
$\lambda(\mathcal{N}_{\text{nz}};t)=-\Delta\dot{m}_{\text{f,c}}(t)=-\Delta\left(\dot{m}_{1,2}(t)-\dot{m}_{2,3}(t)\right)=0\quad\Rightarrow\quad\dot{m}_{1,2}(t)=\dot{m}_{2,3}(t),$ (17)
@@
Currently, CiRL considers microalgae cultivations as the remover compartment (Udaypal et al., 2024). Microalgae are promising for carbon removal because they are a renewable source and, at the end of their life, they can be used as biomass for biofuel production (Udaypal et al., 2024; Sadvakasova et al., 2023). They also have shown to absorb more $\text{CO}_{2}$ than terrestrial plants (Zhou et al., 2017; Wang et al., 2008). The RL controllers currently designed in CiRL aim to improve the ability of microalgae cultivations to absorb $\text{CO}_{2}$ by regulating the light intensity. Hence, the action space is one-dimensional.
@@ 5.2.1 Monod’s Model for Microalgae
One of the simplest models used to describe the dynamics of microalgae cultivations is the Monod’s model (Bernard et al., 2016; Zocco et al., 2025a). It is a state-space model of two states where the growth limiting factors enter the system as multiplying terms in the equation of the growth rate $\mu$. The model implemented in CiRL is detailed in (Vatcheva et al., 2006; Bernard, 2011), while the design of the RL controllers is detailed in (Zocco et al., 2025a). For this paper, we executed a single training with each SB3 algorithm whose performance is reported in Table 6.
Table 6: Training time, $r_{\text{s}}$, $r_{\text{e}}$, and $\zeta$ for each algorithm with CO2MicroalgaeMonod. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$
A2C 08:10 127.7 143.1 15.4
ARS 01:28 114.6 142.8 28.2
@@ 5.2.2 Droop’s Model for Microalgae
The Droop’s model can capture the dynamics of microalgae growth more accurately than the Monod’s model. With respect to the latter, it has an extra state, which models the internal cell quota (Vatcheva et al., 2006; Bernard, 2011). The Droop’s model derives from the application of the mass conservation principle (Saccardo et al., 2023), and hence, is a thermodynamic compartment of a TMN (Zocco et al., 2024). The reward function we implemented for the Droop’s model is that of the Monod’s model, that is (Zocco et al., 2025a),
@@
Table 7: Training time, $r_{\text{s}}$, $r_{\text{e}}$, and $\zeta$ for each algorithm with CO2MicroalgaeDroop. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$
A2C 08:26 409.2 473.0 63.8
@@
SAC 44:17 343.3 459.6 116.3

### v2

@@ 5 Compartments and Compartmental Networks
Reacher $c^{2}_{2,2}$ in Fig. 1 Same as in Gymnasium A2C, ARS, DDPG, PPO, SAC 5.1.1
@@
CarboNet 15 comp. Our implementation in state-space form A2C, ARS, DDPG, PPO, SAC 5.3.1
CopperNet 17 comp. Our implementation in state-space form A2C, ARS, DDPG, PPO, SAC 5.3.2
@@ 5.1 Compartments for Circularity of Solids
The success of sorting $s$ affects the magnitude of $\lambda$. For example, $s$ = 0 yields that all the waste material enters the incinerator at $t_{3,\text{in},6}$, i.e., $\lambda_{\text{a}}$ = $\lambda(\mathcal{N}_{\text{s}};t_{3,\text{in},6},0,T_{\text{s}})$ = 2m-2m. In contrast, $s$ = 100 increases circularity by extending the life-cycle of the material, i.e., $\lambda(\mathcal{N}_{\text{s}};t_{3,\text{in},6},100,T_{\text{s}})$ = $m>\lambda_{\text{a}}$. While all the material is inside the incinerator for $t\geq t_{3,\text{in},5}$ regardless of $s$, i.e.,
@@
the recycled mass $m_{\text{r}}$ has extended the life-cycle of the amount $\nu$ = $t_{3,\text{in},5}$ - $t_{3,\text{in},6}$ if $s>0$, which yields an increase of $m_{\text{r}}$ in $\lambda$ for $t_{3,\text{in},6}\leq t<t_{3,\text{in},5}$ ($\lambda$ is independent of $s$ for $t<t_{3,\text{in},6}$ and for $t\geq t_{3,\text{in},5}$). Extending the life-cycle of materials is a key principle of circular economy (potting2017circular).
@@ 5.1.1 Reacher
The Reacher environment (towers2024gymnasium) consists of a robot with two revolute joints rotating around a direction normal to the plane, it has two degrees of freedom, its actions are the torques applied to the revolute joints, and the goal of the robot is to reach a target point generated at a random location. It is a simplification of the pick-and-place task needed to perform waste sorting. Its simplicity enable us to illustrate how the percentage success $s$ is affected by the RL controller performance, which in turns affects $\lambda$ as shown in Fig. 2. We tested the five SB3 algorithms reported in Table 2, where
@@
with $r_{\text{e}}$ and $r_{\text{s}}$ denoting the mean reward over 100 episodes at the end and at the start of the training, respectively , and where the sample efficiency is quantified via
$\psi=\frac{\zeta}{n_{\text{ts}}},$ (12)
with $n_{\text{ts}}$ the number of training time steps. Thus, $\psi$ is the average increase of the reward per time step, i.e., per interaction between the agent and the environment. The parameter $s$ is calculated by considering as a success an episode in which the end effector reaches a distance from the target smaller than 0.04 m and the joint torques are less than 0.005 Nm. We have run 10,000 episodes and counted the successful episodes.
Table 2: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$, $\psi$, and $s$ for each algorithm with Reacher. Trainings were executed for 100,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$ $s$
A2C 04:37 -63.4 -16.6 46.8 $4.7\times 10^{-4}$ 17.8
ARS 01:00 -12.1 -11.0 1.1 $1.1\times 10^{-5}$ 2.8
DDPG 13:43 -60.2 -5.5 54.7 $5.5\times 10^{-4}$ 72.3
PPO 03:59 -60.2 -37.3 22.9 $2.3\times 10^{-4}$ 20.6
SAC 22:23 -45.1 -8.3 36.8 $3.7\times 10^{-4}$ 51.7
@@ 5.1.2 Fetch Waste Sorting
The Reacher environment is a very simplified model of robotic waste sorting. A more realistic simulator is the one based on the Gymnasium Robotics environment Fetch-PickAndPlace, which is a 7-DOF manipulator that performs pick-and-place with a two-finger gripper (plappert2018multi; towers2024gymnasium). The RL agent controls the opening and closure of the fingers along with the motion of the robot joints. To resemble a waste sorting tasks we modified the Gymnasium environment by constraining the target to be on the table instead of in the air and by locating the target at a fixed point instead of being randomly generated since the coordinates of the point where to place a waste item are usually known and fixed, e.g., the position of the bin. Hence, this environment is referred to as FetchWasteSorting. With this environment, we tested the SB3 algorithms whose performance are reported in Table 3 and in Figs. 5a-5d.
Table 3: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\psi$, and $\zeta$ for each algorithm with FetchWasteSorting. Trainings were executed for 150,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
DDPG-HER 50:06 -50.25 -50 0.25 $1.6\times 10^{-6}$
SAC-HER 63:43 -50.25 -12.67 37.58 $2.5\times 10^{-4}$
TD3-HER 45:12 -50.25 -50 0.25 $1.6\times 10^{-6}$
TQC-HER 65:19 -50.25 -11.45 38.80 $2.6\times 10^{-4}$
@@ 5.1.3 Transport Truck
The transport truck (compartment $c^{6}_{2,3}$) moves the unsorted waste of mass $m_{\text{u}}$ from the sorting facility ($c^{2}_{2,2}$) to the incinerator ($c^{3}_{3,3}$) as shown in Fig. 1. Currently, the model of the truck in CiRL consists of a particle of mass $m_{\text{tot}}=m_{\text{truck}}+m_{\text{u}}$, where $m_{\text{truck}}$ is the mass of the empty truck. As demonstrated by zocco2023thermodynamical, the first law of thermodynamics leads to the Lagrange’s equations, which in turns yield the second Newton’s law of motion (zocco2024circular). Considering that the particle-truck is subject to the traction force $F$, the state-space form of the system is
@@
and hence, the RL controller trained successfully generates a force $F$ that brings the truck to the incinerator located at $x_{\text{inc}}$ and stops it there. This condition corresponds to the maximum reward $r_{\text{max}}=0$. The performance of a single run of the tested RL algorithms is reported in Table 4. It is worth noticing that the proposed model is simplified and does not consider elements that could be important in practice such as the friction, the aerodynamic drag, or real-world terrain variability. For transportation systems based on railway vehicles, for example, the model presented by malvezzi2013tool includes different motion resistance sources.
Table 4: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$, and $\psi$ for each algorithm with TransportTruck. Trainings were executed for 400,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
A2C 18:10 $1.0\times 10^{11}$ $1.2\times 10^{8}$ $\bm{1.0\times 10^{11}}$ $2.5\times 10^{5}$
ARS 02:56 $4.0\times 10^{8}$ $7.1\times 10^{8}$ $3.1\times 10^{8}$
DDPG 50:42 $1.7\times 10^{11}$ $2.0\times 10^{11}$ $3.3\times 10^{10}$
PPO 14:34 $1.03\times 10^{11}$ $1.02\times 10^{11}$ $\bm{8.6\times 10^{8}}$ $2.1\times 10^{3}$
SAC 1:28:18 $9.8\times 10^{10}$ $1.3\times 10^{11}$ $3.4\times 10^{10}$
@@ 5.1.4 Incinerator
The compartment $c^{3}_{3,3}$ in Fig. 1 is the incinerator, which processes end-of-life products and materials. From a circular economy perspective, it is among the least circular options because it destroys the functionality of products and their components (potting2017circular). A complete dynamical model of an incinerator was developed by magnanelli2020dynamic as a result of mass and energy balances. In this first version of CiRL, we consider only the wastebed and the freeboard sub-compartments of magnanelli2020dynamic. This yields the following system of six ordinary differential equations in the state-space form and with state vector $\bm{x}=[M,M_{\text{char}},T_{\text{w}},M_{\text{g,w}},M_{\text{g,f}},T_{\text{g}}]^{\top}=[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}]^{\top}$, where $M$ is the mass of waste in the wastebed, $M_{\text{char}}$ is the mass of char in the wastebed, $T_{\text{w}}$ is the temperature in the wastebed, $M_{\text{g,w}}$ is the mass of the gasses in the wastebed, $M_{\text{g,f}}$ is the mass of the gasses in the freeboard, and $T_{\text{g}}$ is the temperature of the gasses in the freeboard.
@@
The performance of a single run of the tested RL algorithms on the incinerator environment is reported in Table 5 and in Figs. 5e-5h.
Table 5: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$, and $\psi$ for each algorithm with Incinerator. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
A2C 08:25 $3.40\times 10^{5}$ $3.79\times 10^{5}$ $3.9\times 10^{4}$
ARS 01:28 $3.65\times 10^{5}$ $3.53\times 10^{5}$ $\bm{1.21\times 10^{4}}$ $6.0\times 10^{-2}$
DDPG 24:50 $3.23\times 10^{5}$ $3.44\times 10^{5}$ $2.07\times 10^{4}$
PPO 07:30 $3.51\times 10^{5}$ $3.58\times 10^{5}$ $0.71\times 10^{4}$
SAC 43:20 $3.37\times 10^{5}$ $3.41\times 10^{5}$ $0.40\times 10^{4}$
@@ 5.2 Compartments for Circularity of $\text{CO}_{2}$
In this section, we address the circulation of atmospheric carbon dioxide since its accumulation is one of the main causes of climate change (NASA-ClimateChange). Governments usually refer to the need of carbon dioxide circulation as the “net-zero target”, which is defined by NZ-DefUniOx, e.g., the EU (EU-NZ), the UK (UK-NZ), the US (US-NZ), China (China-NZ), and India (India-NZ). Using the TMN formalism, the net-zero problem can be represented as
@@
$\begin{gathered}\lambda(\mathcal{N}_{\text{nz}};t)=-\Delta\dot{m}_{\text{f,c}}(t)=-\Delta\left(\dot{m}_{1,2}(t)-\dot{m}_{2,3}(t)\right)=0\\ \Rightarrow\quad\dot{m}_{1,2}(t)=\dot{m}_{2,3}(t),\end{gathered}$ (18)
@@
Currently, CiRL considers microalgae cultivations as the remover compartment (goswami2024advances). Microalgae are promising for carbon removal because they are a renewable source and, at the end of their life, they can be used as biomass for biofuel production (goswami2024advances; sadvakasova2023microalgae). They also have shown to absorb more $\text{CO}_{2}$ than terrestrial plants (zhou2017bio; wang2008co). The RL controllers currently designed in CiRL aim to improve the ability of microalgae cultivations to absorb $\text{CO}_{2}$ by regulating the light intensity. Hence, the action space is one-dimensional.
@@ 5.2.1 Monod’s Model for Microalgae
One of the simplest models used to describe the dynamics of microalgae cultivations is the Monod’s model (bernard2016modelling; zocco2025circular). It is a state-space model of two states where the growth limiting factors enter the system as multiplying terms in the equation of the growth rate $\mu$. The model implemented in CiRL is detailed in (vatcheva2006experiment; bernard2011hurdles), while the design of the RL controllers is detailed in (zocco2025circular). For this paper, we executed a single training with each SB3 algorithm whose performance is reported in Table 6 and in Figs. 5i-5l.
Table 6: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$, and $\psi$ for each algorithm with CO2MicroalgaeMonod. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
A2C 08:10 127.7 143.1 15.4 $7.7\times 10^{-5}$
ARS 01:28 114.6 142.8 28.2 $1.4\times 10^{-4}$
@@ 5.2.2 Droop’s Model for Microalgae
The Droop’s model can capture the dynamics of microalgae growth more accurately than the Monod’s model. With respect to the latter, it has an extra state, which models the internal cell quota (vatcheva2006experiment; bernard2011hurdles). The Droop’s model derives from the application of the mass conservation principle (saccardo2023droop), and hence, is a thermodynamic compartment of a TMN (zocco2024unification). The reward function we implemented for the Droop’s model is that of the Monod’s model, that is (zocco2025circular),
@@
Table 7: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$ and $\psi$ for each algorithm with CO2MicroalgaeDroop. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
A2C 08:26 409.2 473.0 63.8 $3.2\times 10^{-4}$
@@
SAC 44:17 343.3 459.6 116.3 $5.8\times 10^{-4}$
5.3 Compartmental Networks
This section covers two networks of compartments, namely, CarboNet and CopperNet. The former focuses on the flow of tropospheric carbon dioxide, while the latter on the supply-recovery chain of copper.
5.3.1 CarboNet
CarboNet is a network with 15 compartments and it addresses the circularity of carbon dioxide. It was recently proposed in zocco2025carbonet and controlled via linear quadratic regulators. Here, it is implemented as an SB3 environment to approach its control via reinforcement learning. We set the reward function for this environment as
$r_{\text{ca}}=-(x_{1,\text{d}}-x_{1})^{2},$ (21)
where $x_{1,\text{d}}$ is the desired mass of $\text{CO}_{2}$ in the troposphere of a target area corresponding to the concentration of $\text{CO}_{2}$ in the pre-industrial era. The performance of a single run of the tested RL algorithms on CarboNet is reported in Table 8 and in Figs. 5m-5p.
Table 8: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$ and $\psi$ for each algorithm with CarboNet. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
A2C 07:36 $7.70\times 10^{6}$ $7.06\times 10^{7}$ $6.29\times 10^{7}$
ARS 01:24 $3.47\times 10^{7}$ $2.04\times 10^{6}$ $\bm{3.27\times 10^{7}}$ 163
DDPG 24:07 $4.10\times 10^{7}$ $7.08\times 10^{7}$ $2.97\times 10^{7}$
PPO 06:43 $13\times 10^{7} 1.1$$17\times 10^{7}$ $3.70\times 10^{5}$
SAC 39:27 $2.66\times 10^{7}$ $7.08\times 10^{7}$ $4.41\times 10^{7}$
5.3.2 CopperNet
This environment, namely, CopperNet, is made of six linear ordinary differential equations resulting from the mass balance principle applied to six nodes of a supply-recovery chain that extracts, transports, transforms, and use copper. The network architecture is inspired by the material flow analysis in (loibl2021current, Fig. 1). The compartmental diagraph of CopperNet is shown in Fig. 4,
Figure 4: Compartmental digraph of a supply-recovery network of copper, namely, CopperNet.
while its state-space model is given by
$\dot{\bm{x}}=\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ \dot{x}_{3}\\ \dot{x}_{4}\\ \dot{x}_{5}\\ \dot{x}_{6}\end{bmatrix}=\begin{bmatrix}a_{17}x_{1}+a_{16}x_{6}-a_{21}x_{1}\\ a_{21}x_{1}+a_{26}x_{6}-a_{32}x_{2}\\ a_{32}x_{2}-a_{43}x_{3}\\ a_{43}x_{3}-a_{84}x_{4}-u\\ -a_{65}x_{5}+u\\ a_{65}x_{5}-a_{16}x_{6}-a_{26}x_{6}\end{bmatrix}$, (22)
where $a_{i,j}\geq$ 0 [1/day] is the rate constant that is proportional to the mass flow rate between the node $i$ and the node $j$ of the digraph (see Fig. 4). The six states $x_{i}$ [kt], $i=1,\dots,6$, are masses located at different stages of the supply-recovery chain of copper as indicated in Fig. 4, while $u$ [kt/day] is the action of the RL agent. Coherently with the notation in Definition 2, CopperNet can be compactly represented by the set
$\begin{gathered}\mathcal{N}_{\text{c}}=\{c^{1}_{1,1},c^{2}_{2,2},c^{3}_{3,3},c^{4}_{4,4},c^{5}_{5,5},c^{6}_{6,6},c^{7}_{7,7},c^{8}_{8,8},c^{9}_{1,2},\\ c^{10}_{2,3},c^{11}_{3,4},c^{12}_{4,5},c^{13}_{4,8},c^{14}_{5,6},c^{15}_{6,1},c^{16}_{6,2},c^{17}_{7,1}\}.\end{gathered}$ (23)
Thus, for $\mathcal{N}_{\text{c}}$ it holds that $n_{\text{c}}$ = 17, $n_{\text{v}}$ = 8, and $n_{\text{a}}$ = 9. By applying Definition 4, the circularity of CopperNet is
$\lambda=-(a_{17}x_{1}+a_{84}x_{4}),$ (24)
where the first term on the right-hand side of the equation takes into account the finite-time sustainable extraction of copper, whereas the second term on the right-hand side of the equation takes into account the leaks of copper that end-up in landfills, incinerators, and the natural environment.
Finally, we defined the reward function as
$r_{\text{co}}=-\sum_{i=1}^{6}(x_{i,\text{d}}-x_{i})^{2},$ (25)
where $x_{i,\text{d}}$ is the desired value of the state $x_{i}$. The performance of a single run of the tested RL algorithms on CopperNet is reported in Table 9 and in Figs. 5q-5t.
Table 9: Training time, $r_{\text{s}}$, $r_{\text{e}}$, $\zeta$, and $\psi$ for each algorithm with CopperNet. Trainings were executed for 200,000 time steps on a Colab NVIDIA Tesla T4. In bold values of $\zeta>0$. Sample efficiency $\psi$ indicated only for $\zeta>0$.
Tr. time (min:sec) $r_{\text{s}}$ $r_{\text{e}}$ $\zeta$ $\psi$
A2C 07:26 $1.34\times 10^{8}$ $5.30\times 10^{7}$ $\bm{8.15\times 10^{7}}$ 407
ARS 01:28 $2.81\times 10^{8}$ $4.71\times 10^{7}$ $\bm{2.34\times 10^{8}}$ 1170
DDPG 22:56 $2.57\times 10^{8}$ $4.43\times 10^{8}$ $1.86\times 10^{8}$
PPO 06:38 $9.32\times 10^{7}$ $8.50\times 10^{7}$ $\bm{0.82\times 10^{7}}$ 41
SAC 41:25 $2.67\times 10^{8}$ $4.43\times 10^{8}$ $1.76\times 10^{8}$
(a) DDPG-HER
(b) SAC-HER
(c) TD3-HER
(d) TQC-HER
(e) A2C
(f) ARS
(g) DDPG
(h) PPO
(i) A2C
(j) ARS
(k) PPO
(l) SAC
(m) A2C
(n) ARS
(o) DDPG
(p) PPO
(q) A2C
(r) ARS
(s) DDPG
(t) PPO
Figure 5: Learning curves of different algorithms with the following environments: (a)-(d) FetchWasteSorting; (e)-(h) Incinerator; (i)-(l) CO2MicroalgaeMonod; (m)-(p) CarboNet; (q)-(t) CopperNet. The learning curves yield a qualitative evaluation of the convergence stability (rahim2024tinyfdrl).